OpenStudy (anonymous):

PLEASE SOLVE THIS: Solve the equation 3p + 2 = 5 + 3p

OpenStudy (cwrw238):

subtract 3p from both sides: what do you get?

OpenStudy (cwrw238):

2 = 5 right?

OpenStudy (anonymous):

you will want to subtract the 2 form the 5 giving you 3p=5+3p then subtract the 3p and what does tht give you?

OpenStudy (anonymous):

3p=2*+3p

OpenStudy (cwrw238):

so the solution is there is no solution

OpenStudy (anonymous):

@cwrw238 the solution is 2

OpenStudy (anonymous):

I believe there is no solution because p equals zero

OpenStudy (anonymous):

Is that one of your choices

OpenStudy (cwrw238):

how can the solution be 2? plug p=2 into the equation 3(2) + 2 = 8 5 + 3(2)

OpenStudy (cwrw238):

5 + 3(2) = 11

OpenStudy (cwrw238):

there is no value of p which will satisfy this equation